# Evaluating a VR System for Collecting Safety-Critical Vehicle-Pedestrian Interactions

### v1

@@ Abstract.
*Corresponding author
**work done while visiting at Carnegie Mellon University
submissionid: 8406
@@
Figure 1. Left: Participant’s avatar within the virtual environment from a third-person perspective. Right: participant wearing the VR headset in the data collection environment.
1. Introduction
@@
Pedestrian trajectory data collection is necessary for training safe and robust pedestrian forecasting models for autonomous vehicles (AVs), but many important vehicle-pedestrian interaction scenarios are missing in existing popular trajectory and tracking datasets (Caesar et al., 2019; Ettinger et al., 2021; Geiger et al., 2013; Chang et al., 2019; Wilson et al., 2023). Specifically, collecting data in many scenarios is difficult. For example, it is difficult to collect data of jaywalkers: pedestrians who cross a street at a location other than designated crosswalks or intersections, often in disregard of traffic regulations. Jaywalking is generally considered unsafe and is illegal in many regions, and thus it is sparsely observed in real life or real datasets. However, it is still important that forecasting models understand the behavior of jaywalkers for pedestrian safety. Furthermore, it is problematic to willfully direct pedestrians to perform illegal behavior for the purpose of data collection. As another example, real-life datasets have a deficiency of data of children pedestrians (Caesar et al., 2019; Ettinger et al., 2021; Geiger et al., 2013; Chang et al., 2019; Wilson et al., 2023). Child street-crossing behavior differs from that of adults; children often take more risks and appear more unpredictable (Gitelman et al., 2019). Due to child protection laws, it is even more problematic to put children in potentially dangerous situations for the purpose of data collection (Office for Human Research Protections (2010), OHRP). As a final example, there exist very specific vehicle-pedestrian interaction scenarios, such as walking on the road alongside traffic, crossing a parking lot or garage entrance with vehicles entering and exiting, or crossing a narrow urban alleyway in which the building structure obstruct easy view of oncoming vehicles. These scenarios arise due to specific urban structural and geographic features that may only be sparsely found in real-life urban areas. However, these scenarios are still relevant and important to robust pedestrian safety, and thus it is important that data is collected for these scenarios. There is a large body of work on controllable simulators that can generate large quantities of synthetic data of various scenarios of choice (Dosovitskiy et al., 2017; Chen et al., 2021; Yang et al., 2023), including the aforementioned scenarios that are uncommon in real-life but critical to pedestrian safety. There is also much work done on AV perception, prediction, and planning models that utilizes simulated data for training (Johnson-Roberson et al., 2016; Huang and Ramanan, 2017a). Relevant to our area of interest, Huang and Ramanan (2017b) utilize synthetic pedestrian data to train a detector to detect unusual or uncommon pedestrian behaviors. However, all such methods that train largely on synthetic data suffer from the well-known ”sim2real” gap. This gap exists because simulators struggle to reproduce certain aspects of reality: for example, in CARLA (Dosovitskiy et al., 2017), vehicles travel along “rails” located in the center of the lane, never deviating from the center as real human drivers do. In addition, CARLA’s pedestrian models move mechanically according to mathematical gait functions. Due to the complexity of human behavior and movement characteristics, pedestrians are particularly difficult to simulate.
@@
These limitations point to the need for other solutions to obtain such data. Virtual Reality (VR) offers a solution for collecting pedestrian data in a safe environment free of real automotive dangers; VR can be used to immerse a pedestrian in a virtual scenario analogous to one found in real-life, and the pedestrian can be asked to cross a virtual road, or traverse the virtual environment, all the while their trajectory data is collected. In addition, data collected from VR is collected from real humans, and thus does not suffer from the same human behavioral sim2real gap as simulated data. However, VR may exhibit limitations to realistic data collection. For example, the VR user experience could be unrealistic, whether it’s because of the graphics of the virtual environment, lag or delay in the VR view refresh frame rate, or the weight of the VR headset on their head. These issues may interfere with the pedestrian’s sense of immersion, and accordingly, may alter their actions and movements from what they would do on a real street. Evaluating a VR system (Anonymized, 2023) that was developed concurrently with this work through empirical studies, we show that although users do not experience perfect sense of presence or immersion within the system, data collected from the system is mostly indistinguishable from that collected in real life, providing evidence for our claim that the VR system is effective at collecting a genuine pedestrian response to vehicle motion. Our contributions are as follows:
  1. (1)
@@
2. Background and Related Work
2.1. Human Behavior in Virtual Environments
@@
Many works have studied human behavior within virtual environments. Some works study the efficacy and quality of the VR experience via objective measures such as stride length and gait (Janeh et al., 2018) as well as subjective measures such as subjects’ ratings on presence questionnaires  (Drey et al., 2021; Berkman and Akan, 2019; Slater and Wilbur, 1997; Stanney et al., 2003). Some study specific behaviors of interest like normal walking (Nilsson et al., 2018), evacuating a building (Arias et al., 2022), collision-avoidance  (Berton et al., 2019), proxemics and group behavior (Podkosova and Kaufmann, 2018; Norouzi et al., 2019), etc. Some focus on behaviors in specific professional areas like surgery (McCloy and Stone, 2001; Li et al., 2017) or education (Bell and Scott Fogler, [n. d.]). Many methods evaluate the efficacy of VR systems by comparing human behavior in VR vs. real environments (Merino et al., 2020). Finally, other works evaluate how different conditions of the VR system’s features affect user sense of presence behavior, such as level of photorealism of the environment rendering (Zibrek and McDonnell, 2019), locomotion methods (Tan et al., 2022; Nilsson et al., 2015; Olivier et al., 2018), and avatar appearance (Samaraweera et al., 2013; Ogawa et al., 2020).
2.2. Pedestrian Behavior Studies
@@
Extensive study has been done to characterize pedestrian behavior in traffic environments and responses to vehicle behaviors (Jayaraman et al., 2018; Colley and Rukzio, 2020; Lee et al., 2019; Feng et al., 2021; Weber et al., 2019; Chen et al., 2020). Some works study AV behavior or evaluate improvements to the AV-pedestrian communication interface such as intent signalling methods (Jayaraman et al., 2018; Colley and Rukzio, 2020; Lee et al., 2019; Feng et al., 2021; Lee et al., 2019; Fuest et al., 2020; Ackermans et al., 2020; Deb et al., 2020; de Clercq et al., 2019). Many of the aforementioned works, as well as others, use immersive VR environments to simulate scenarios of interest and study pedestrian’s responses (Nuñez Velasco et al., 2019; Schneider and Bengler, 2020; Bhagavathula et al., 2018, 2018; Feldstein and Dyszak, 2020; Mahadevan et al., 2019; Tran et al., 2021; Nuñez Velasco et al., 2019). Some works focus on evaluating the efficacy of VR systems for use in studying human behavior or evaluating safety systems or vehicle-pedestrian communication systems by comparing subject’s responses in virtual environments to that of real-world environments (Nascimento et al., 2019; Deb et al., 2017; Schneider and Bengler, 2020; Higuera-Trujillo et al., 2017). Work on VR-based dataset collection and benchmarking tools similar to that studied in this work has also been done, such as (Dalipi et al., 2020), but their focus is on testing vehicle-pedestrian communication interfaces rather than on collecting realistic vehicle-pedestrian interaction behavior.
@@
2.3. 2D Pedestrian Trajectory Data Capture: Real-Life, Human-Informed, and Fully-Synthetic
@@
Previous works have used various methods to capture vehicle-pedestrian interaction datasets. The most popular datasets, naturally, are those collected in real-life environments  (Pellegrini et al., 2009; Lerner et al., 2007; Robicquet et al., 2016). PedX (Kim et al., 2018) and PIE (Rasouli et al., 2019) contain pedestrian trajectories collected at intersections, which is the closest data to our use case. However, they too are limited in scope; PedX contains full body pose data but in limited environments such as crosswalks, and PIE contains only 2D bounding box annotations. Popular open-source vehicle datasets also contain pedestrians (Chang et al., 2019; Wilson et al., 2023; Caesar2019). However, as verified in the first part of our study in Section 3, these datasets lack diverse pedestrian behavior and diverse vehicle-pedestrian interactions. Most vehicle-pedestrian interactions in these datasets are at intersection crosswalks.
@@
Some datasets are completely synthetic, collected autonomously via simulated environments that operate based on rules. While some simulators only produce simplistic point trajectory behavior (Biswas et al., 2021; Kothari et al., 2022), others are extensively customizable and controllable worlds that can model complex sensor data (Dosovitskiy et al., 2017; Chen et al., 2021; Yang et al., 2023). Although simulators simplify the process of collecting large-scale data in diverse environments, there is a domain gap between synthetic and real data that is often difficult to measure, and many simulators do not reproduce human behavior well. Because AV models ultimately must work for real pedestrians and vehicles, in this way synthetic data falls short.
@@
Some methods, such as driving simulators (Silvera et al., 2022; Michalík et al., 2021), invite human input to elicit real human behavior. The Garden of Forking Paths (Liang et al., 2019) collects annotations of trajectory continuations from human annotators to create a multi-future pedestrian trajectory dataset in simulated environments reconstructed from real-world scenes. However, they only annotate trajectories at the 2D point-level, through a very limited keyboard and mouse interface, a far cry from the controllability of a VR headset and environment.
3. Part 1: Understanding Limitations of Existing Vehicle-Pedestrian Datasets
@@
3.1. Semi-Structured Interview Recruitment and Design
@@
3.2. Interview Responses
To analyze the state of the limitations of current trajectory datasets, we used inductive thematic analysis (Braun and Clarke, 2006), grouping the themes that arose during the interviews into four distinct categories, which we describe below.
3.2.1. Lack of Uncommon but Important Scenarios
@@
3.2.2. Lack of Interesting Vehicle-Pedestrian Interactions
Both I1 and I2 pointed out that the most popular public vehicle-pedestrian datasets lack “interesting interactions.” One of the most popular datasets (Chang et al., 2019; Wilson et al., 2023) in particular contains a subset that was compiled specifically for the “interestingness” of its scenarios based on internally-defined heuristics, such as number of other agents present in the scene, speed changes, and lane changes. However, I1 remarked that these scenarios were still mostly uninteresting, claiming that it is difficult to come up with a straightforward heuristic to separate ”interesting” scenes from ”boring” scenes. I2 stated that around 75-80% of the data they worked with is “uneventful”, and that there was insufficient diversity in scenes. Scenes often featured empty or minimally populated roads where only the ego-vehicle, the vehicle performing the data collection, was present. Vehicles also lacked varied behaviors such as lane changes. I3 claimed that the pedestrian-only datasets they worked with also lacked variety in environment layouts and did not contain many pedestrian interactions.
3.2.3. Lack of Fine-grained Trajectory Features
Vehicle-pedestrian datasets lack richness in representation. I1 commented that existing trajectory datasets are heavily preprocessed from raw lidar data, losing a lot of the richness of the raw representation. Popular trajectory datasets such as Argoverse (Chang et al., 2019; Wilson et al., 2023) and NuScenes (Caesar et al., 2019) represent pedestrian trajectories as 2D points rather than full 3D human bodies. I2 notes that while pedestrian body orientation is also recorded, it still falls far from the detail available in full body pose. In tracking datasets, agents are often annotated with simple 3D rectangular bounding boxes. I3 explained that pedestrian bounding boxes often overlap with one another, because humans occupy their bounding boxes only sparsely. Overlapping bounding boxes appear to be colliding; this may lead to model training issues such as failure to understand collision-avoidance. Thus, bounding boxes also fall short in representing the richness of pedestrian motion.
@@ 3.2.4. Summary
4. VR Pedestrian Trajectory and Pose Collection System Overview
@@
Figure 2. Schematic of the VR data collection system.
Figure 3. Left: The VR system set up in the classroom data collection space. Right: A pedestrian subject wearing the VR headset.
Because the engineering and design of the VR pedestrian simulator is not the primary contribution of this paper but of (Anonymized, 2023), we leave further details about the data collection system to that work.
5. Part 2: VR Sense of Presence User Study
@@ 5.1. Study Design
Figure 4. The three tasks that we ask our study participants to complete. Top: snapshot of each task as seen in the CARLA virtual environment. Bottom: schematic of each task; “S” denotes participant start location and “G” denotes goal. Schematic images also used in Anonymized (2023).
5.2. Self-Reported Measures
@@
To design the questions asked in the post-experience questionnaire, we used a combination of questions derived from previous works, questions modified from previous works, and custom questions designed specifically to evoke defining attributes of the VR user experience. We designed 12 quantitative questions using a semantic differential scale (Tina Øvad, [n. d.]) from 1 to 7, as well as 4 free-response questions with free text entry. We grouped the questions into 3 categories reflecting different aspects of trustworthiness of the data collection system: sense of presence experienced within the virtual environment (Berkman and Akan, 2019; Witmer and Singer, 1998) (coded by the letter P), sense of agency (A), and behavioral and experiential similarity to real-life (B). In total, the post-condition questionnaire was composed of 16 questions aimed to cover a range of subjective ratings while keeping the time for participants to complete the questionnaire within 10 minutes. Questions, sources, categorizations, and semantic differential anchors for the questions are recorded in Table 1.
@@
Participants also reported various demographics: age, gender, VR experience (5-point Likert-type scale), video game frequency (5-point Likert-type scale), frequency of jaywalking behavior (5-point Likert-type scale), and level of alertness (Stanford Sleepiness Scale (Shahid et al., 2012)). Summary of participant demographics is reported in Table 2.
ID Source Category Question Semantic Differential Anchors (1normal-$\rightarrow$ 7)
P1
SUS (Usoh et al., 2000)
presence
@@
Please rate your sense of “being there” in the environment.
did not have a sense of ”being there” $\rightarrow$ normal experience of “being there”
P2
SUS (Usoh et al., 2000)
presence
@@
When you think back on your experience, do you think of the environment more as images that you saw, or more as somewhere that you visited?
Images that I saw $\rightarrow$ somewhere that I visited
P3
WS (Witmer and Singer, 1998)
presence
How real did the objects in the environment seem?
@@
very fake, clearly images $\rightarrow$ very real, like I could touch them
P4
custom
presence
@@
During the VR experience, were you more concerned with the real world (this classroom) or the virtual world?
real world $\rightarrow$ virtual world
P5
SUS (Usoh et al., 2000)
presence
@@
When you think back on your experience, do you think of the vehicles more as images that you saw, or as things you interacted with?
Images that I saw $\rightarrow$ things that I interacted with
P6
custom
presence
@@
During the experience, did you often think to yourself that the vehicles were physical objects that could have actually hit you and caused you injury?
very much so $\rightarrow$ not at all
A1
custom
agency
@@
How comfortable did you feel moving around in the environment?
very uncomfortable $\rightarrow$ very comfortable
A2
AE (Gonzalez-Franco and Peck, 2018)
agency
How much did you feel like you could control the virtual body?
@@
did not feel much agency $\rightarrow$ could control it like own body
A3
SPES (Hartmann et al., 2016)
agency
@@
How freely did you feel you could move in the environment?
movements were restricted $\rightarrow$ movements were free
B3
NASA-TLX (Hart and Staveland, 1988)
behavior
@@
How mentally demanding were the tasks compared to doing them in the real world (e.g. on a real street)?
less demanding $\rightarrow$ more demanding
B1
custom
behavior
@@
Did you feel your head, arm, and body movements were the same as they would have been in the real world?
completely different $\rightarrow$ exactly identical
BF1
custom
behavior
@@
What parts of your movements were different than how they would have been in the real world?
free response
B2
custom
behavior
@@
Did you feel your decisions about when to act were the same as they would have been in the real world?
completely different $\rightarrow$ exactly identical
BF2
custom
behavior
@@
What parts of your decisions were different than how they would have been in the real world?
free response
PF1
custom
presence
@@
What aspects of the systems or environment were realistic?
free response
PF2
custom
presence
@@
What aspects of the system or environment were unrealistic?
free response
Table 1. Post-condition questionnaire used in the study.
@@
Level of alertness {0-7} (Shahid et al., 2012) 6 {3-7}
Table 2. Summary of participant demographics. Continuous variables are summarized as mean (standard deviation) and ordinal variables are summarized as median {range}.
5.3. Results and Analyses
Figure 5. Various study participants wearing the VR headset and walking around in two different classroom-setting data collection environments.
5.3.1. Quantitative Measures

### v3

@@ Abstract
Figure 1: Left: Participant’s avatar within the virtual environment from a third-person perspective. Right: participant wearing the VR headset in the data collection environment.
I Introduction
@@
Pedestrian trajectory data collection is necessary for training safe and robust pedestrian forecasting models for autonomous vehicles (AVs), but many important vehicle-pedestrian interaction scenarios are missing in existing popular trajectory and tracking datasets [11, 22, 26, 12, 70]. Specifically, collecting data in many scenarios is difficult. For example, it is difficult to collect data of jaywalkers: pedestrians who cross a street at a location other than designated crosswalks or intersections, often in disregard of traffic regulations. Jaywalking is generally considered unsafe and is illegal in many regions, and thus it is sparsely observed in real life or real datasets. However, it is still important that forecasting models understand the behavior of jaywalkers for pedestrian safety. Furthermore, it is problematic to willfully direct pedestrians to perform illegal behavior for the purpose of data collection. As another example, real-life datasets have a deficiency of data of children pedestrians [11, 22, 26, 12, 70]. Child street-crossing behavior differs from that of adults; children often take more risks and appear more unpredictable [27]. Due to child protection laws, it is even more problematic to put children in potentially dangerous situations for the purpose of data collection [52]. As a final example, there exist very specific vehicle-pedestrian interaction scenarios, such as walking on the road alongside traffic, crossing a parking lot or garage entrance with vehicles entering and exiting, or crossing a narrow urban alleyway in which the building structure obstruct easy view of oncoming vehicles. These scenarios arise due to specific urban structural and geographic features that may only be sparsely found in real-life urban areas. However, these scenarios are still relevant and important to robust pedestrian safety, and thus it is important that data is collected for these scenarios. There is a large body of work on controllable simulators that can generate large quantities of synthetic data of various scenarios of choice [20, 14, 72], including the aforementioned scenarios that are uncommon in real-life but critical to pedestrian safety. There is also much work done on AV perception, prediction, and planning models that utilizes simulated data for training [36, 33]. Relevant to our area of interest, Huang and Ramanan [32] utilize synthetic pedestrian data to train a detector to detect unusual or uncommon pedestrian behaviors. However, all such methods that train largely on synthetic data suffer from the well-known ”sim2real” gap. This gap exists because simulators struggle to reproduce certain aspects of reality: for example, in CARLA [20], vehicles travel along “rails” located in the center of the lane, never deviating from the center as real human drivers do. In addition, CARLA’s pedestrian models move mechanically according to mathematical gait functions. Due to the complexity of human behavior and movement characteristics, pedestrians are particularly difficult to simulate.
@@
These limitations point to the need for other solutions to obtain such data. Virtual Reality (VR) offers a solution for collecting pedestrian data in a safe environment free of real automotive dangers; VR can be used to immerse a pedestrian in a virtual scenario analogous to one found in real-life, and the pedestrian can be asked to cross a virtual road, or traverse the virtual environment, all the while their trajectory data is collected. In addition, data collected from VR is collected from real humans, and thus does not suffer from the same human behavioral sim2real gap as simulated data. However, VR may exhibit limitations to realistic data collection. For example, the VR user experience could be unrealistic, whether it’s because of the graphics of the virtual environment, lag or delay in the VR view refresh frame rate, or the weight of the VR headset on their head. These issues may interfere with the pedestrian’s sense of immersion, and accordingly, may alter their actions and movements from what they would do on a real street. Evaluating a VR system [3] that was developed concurrently with this work through empirical studies, we show that although users do not experience perfect sense of presence or immersion within the system, data collected from the system is mostly indistinguishable from that collected in real life, providing evidence for our claim that the VR system is effective at collecting a genuine pedestrian response to vehicle motion. Our contributions are as follows:
  1. 1.
@@
II Background and Related Work
II-A Human Behavior in Virtual Environments
@@
Many works have studied human behavior within virtual environments. Some works study the efficacy and quality of the VR experience via objective measures such as stride length and gait [34] as well as subjective measures such as subjects’ ratings on presence questionnaires  [21, 6, 63, 64]. Some study specific behaviors of interest like normal walking [49], evacuating a building [4], collision-avoidance  [7], proxemics and group behavior [56, 50], etc. Some focus on behaviors in specific professional areas like surgery [44, 41] or education [5]. Many methods evaluate the efficacy of VR systems by comparing human behavior in VR vs. real environments [45]. Finally, other works evaluate how different conditions of the VR system’s features affect user sense of presence behavior, such as level of photorealism of the environment rendering [73], locomotion methods [65, 48, 54], and avatar appearance [59, 53].
II-B Pedestrian Behavior Studies
@@
Extensive study has been done to characterize pedestrian behavior in traffic environments and responses to vehicle behaviors [35, 15, 39, 24, 69, 13]. Some works study AV behavior or evaluate improvements to the AV-pedestrian communication interface such as intent signalling methods [35, 15, 39, 24, 39, 25, 2, 19, 17]. Many of the aforementioned works, as well as others, use immersive VR environments to simulate scenarios of interest and study pedestrian’s responses [51, 60, 8, 8, 23, 43, 67, 51]. Some works focus on evaluating the efficacy of VR systems for use in studying human behavior or evaluating safety systems or vehicle-pedestrian communication systems by comparing subject’s responses in virtual environments to that of real-world environments [47, 18, 60, 31]. Work on VR-based dataset collection and benchmarking tools similar to that studied in this work has also been done, such as [16], but their focus is on testing vehicle-pedestrian communication interfaces rather than on collecting realistic vehicle-pedestrian interaction behavior.
@@
II-C 2D Pedestrian Trajectory Data Capture: Real-Life, Human-Informed, and Fully-Synthetic
@@
Previous works have used various methods to capture vehicle-pedestrian interaction datasets. The most popular datasets, naturally, are those collected in real-life environments  [55, 40, 58]. PedX [37] and PIE [57] contain pedestrian trajectories collected at intersections, which is the closest data to our use case. However, they too are limited in scope; PedX contains full body pose data but in limited environments such as crosswalks, and PIE contains only 2D bounding box annotations. Popular open-source vehicle datasets also contain pedestrians [12, 70, Caesar2019]. However, as verified in the first part of our study in Section III, these datasets lack diverse pedestrian behavior and diverse vehicle-pedestrian interactions. Most vehicle-pedestrian interactions in these datasets are at intersection crosswalks.
@@
Some datasets are completely synthetic, collected autonomously via simulated environments that operate based on rules. While some simulators only produce simplistic point trajectory behavior [9, 38], others are extensively customizable and controllable worlds that can model complex sensor data [20, 14, 72]. Although simulators simplify the process of collecting large-scale data in diverse environments, there is a domain gap between synthetic and real data that is often difficult to measure, and many simulators do not reproduce human behavior well. Because AV models ultimately must work for real pedestrians and vehicles, in this way synthetic data falls short.
@@
Some methods, such as driving simulators [62, 46], invite human input to elicit real human behavior. The Garden of Forking Paths [42] collects annotations of trajectory continuations from human annotators to create a multi-future pedestrian trajectory dataset in simulated environments reconstructed from real-world scenes. However, they only annotate trajectories at the 2D point-level, through a very limited keyboard and mouse interface, a far cry from the controllability of a VR headset and environment.
III Part 1: Understanding Limitations of Existing Vehicle-Pedestrian Datasets
@@
III-A Semi-Structured Interview Recruitment and Design
@@
III-B Interview Responses
To analyze the state of the limitations of current trajectory datasets, we used inductive thematic analysis [10], grouping the themes that arose during the interviews into four distinct categories, which we describe below.
III-B1 Lack of Uncommon but Important Scenarios
@@
III-B2 Lack of Interesting Vehicle-Pedestrian Interactions
Both I1 and I2 pointed out that the most popular public vehicle-pedestrian datasets lack “interesting interactions.” One of the most popular datasets [12, 70] in particular contains a subset that was compiled specifically for the “interestingness” of its scenarios based on internally-defined heuristics, such as number of other agents present in the scene, speed changes, and lane changes. However, I1 remarked that these scenarios were still mostly uninteresting, claiming that it is difficult to come up with a straightforward heuristic to separate ”interesting” scenes from ”boring” scenes. I2 stated that around 75-80% of the data they worked with is “uneventful”, and that there was insufficient diversity in scenes. Scenes often featured empty or minimally populated roads where only the ego-vehicle, the vehicle performing the data collection, was present. Vehicles also lacked varied behaviors such as lane changes. I3 claimed that the pedestrian-only datasets they worked with also lacked variety in environment layouts and did not contain many pedestrian interactions.
III-B3 Lack of Fine-grained Trajectory Features
Vehicle-pedestrian datasets lack richness in representation. I1 commented that existing trajectory datasets are heavily preprocessed from raw lidar data, losing a lot of the richness of the raw representation. Popular trajectory datasets such as Argoverse [12, 70] and NuScenes [11] represent pedestrian trajectories as 2D points rather than full 3D human bodies. I2 notes that while pedestrian body orientation is also recorded, it still falls far from the detail available in full body pose. In tracking datasets, agents are often annotated with simple 3D rectangular bounding boxes. I3 explained that pedestrian bounding boxes often overlap with one another, because humans occupy their bounding boxes only sparsely. Overlapping bounding boxes appear to be colliding; this may lead to model training issues such as failure to understand collision-avoidance. Thus, bounding boxes also fall short in representing the richness of pedestrian motion.
@@ III-B4 Summary
IV VR Pedestrian Trajectory and Pose Collection System Overview
@@
Figure 2: Schematic of the VR data collection system.
Figure 3: Left: The VR system set up in the classroom data collection space. Right: A pedestrian subject wearing the VR headset.
Because the engineering and design of the VR pedestrian simulator is not the primary contribution of this paper but of [3], we leave further details about the data collection system to that work.
V Part 2: VR Sense of Presence User Study
@@ V-A Study Design
Figure 4: The three tasks that we ask our study participants to complete. Top: snapshot of each task as seen in the CARLA virtual environment. Bottom: schematic of each task; “S” denotes participant start location and “G” denotes goal. Schematic images also used in Anonymized [3].
V-B Self-Reported Measures
@@
To design the questions asked in the post-experience questionnaire, we used a combination of questions derived from previous works, questions modified from previous works, and custom questions designed specifically to evoke defining attributes of the VR user experience. We designed 12 quantitative questions using a semantic differential scale [66] from 1 to 7, as well as 4 free-response questions with free text entry. We grouped the questions into 3 categories reflecting different aspects of trustworthiness of the data collection system: sense of presence experienced within the virtual environment [6, 71] (coded by the letter P), sense of agency (A), and behavioral and experiential similarity to real-life (B). In total, the post-condition questionnaire was composed of 16 questions aimed to cover a range of subjective ratings while keeping the time for participants to complete the questionnaire within 10 minutes. Questions, sources, categorizations, and semantic differential anchors for the questions are recorded in Table I.
@@
Participants also reported various demographics: age, gender, VR experience (5-point Likert-type scale), video game frequency (5-point Likert-type scale), frequency of jaywalking behavior (5-point Likert-type scale), and level of alertness (Stanford Sleepiness Scale [61]). Summary of participant demographics is reported in Table II.
[…]
@@
P1 SUS [68] presence Please rate your sense of “being there” in the environment. did not have a sense of ”being there” $\rightarrow$ normal experience of “being there”
@@
P2 SUS [68] presence When you think back on your experience, do you think of the environment more as images that you saw, or more as somewhere that you visited? Images that I saw $\rightarrow$ somewhere that I visited
@@
P3 WS [71] presence How real did the objects in the environment seem? very fake, clearly images $\rightarrow$ very real, like I could touch them
@@
P4 custom presence During the VR experience, were you more concerned with the real world (this classroom) or the virtual world? real world $\rightarrow$ virtual world
@@
P5 SUS [68] presence When you think back on your experience, do you think of the vehicles more as images that you saw, or as things you interacted with? Images that I saw $\rightarrow$ things that I interacted with
@@
P6 custom presence During the experience, did you often think to yourself that the vehicles were physical objects that could have actually hit you and caused you injury? very much so $\rightarrow$ not at all
@@
A1 custom agency How comfortable did you feel moving around in the environment? very uncomfortable $\rightarrow$ very comfortable
@@
A2 AE [28] agency How much did you feel like you could control the virtual body? did not feel much agency $\rightarrow$ could control it like own body
@@
A3 SPES [30] agency How freely did you feel you could move in the environment? movements were restricted $\rightarrow$ movements were free
@@
B3 NASA-TLX [29] behavior How mentally demanding were the tasks compared to doing them in the real world (e.g. on a real street)? less demanding $\rightarrow$ more demanding
@@
B1 custom behavior Did you feel your head, arm, and body movements were the same as they would have been in the real world? completely different $\rightarrow$ exactly identical
@@
BF1 custom behavior What parts of your movements were different than how they would have been in the real world? free response
@@
B2 custom behavior Did you feel your decisions about when to act were the same as they would have been in the real world? completely different $\rightarrow$ exactly identical
@@
BF2 custom behavior What parts of your decisions were different than how they would have been in the real world? free response
@@
PF1 custom presence What aspects of the systems or environment were realistic? free response
@@
PF2 custom presence What aspects of the system or environment were unrealistic? free response
TABLE I: Post-condition questionnaire used in the study.
@@
Level of alertness {0-7} [61] 6 {3-7}
TABLE II: Summary of participant demographics. Continuous variables are summarized as mean (standard deviation) and ordinal variables are summarized as median {range}.
V-C Results and Analyses
Figure 5: Various study participants wearing the VR headset and walking around in two different classroom-setting data collection environments.
V-C1 Quantitative Measures